OpenStudy (anonymous):

How many sides does the regular polygon have if each interior angle measure is four times the measure of each exterior angle measure.

OpenStudy (ranga):

Assume the polygon has N sides. The exterior angle of a polygon with N sides = 360 / N The interior angle = 180 - exterior angle = 180 - 360 / N "each interior angle measure is four times the measure of each exterior angle measure": 180 - 360 / N = 4 (360 / N) Solve for N

OpenStudy (anonymous):

Is it a double digit number?

OpenStudy (ranga):

Yes. (The lowest double digit number)

OpenStudy (anonymous):

10?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

thank you so much

OpenStudy (ranga):

you are welcome.